OpenStudy (anonymous):

Find the coordinates of the reflection: B (4, 2) is reflected over the line y=1 B’=? Question 9 options: a) (3, 2) b) (-4, 1) c) (4, 0) d) (4, 1)

OpenStudy (anonymous):

use formula p(x,y)= Q ( x, 2k-y) for y=k

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

\[p(x,y) \rightarrow p^l ( x^ly^l) \]

OpenStudy (anonymous):

\[x^l= x \]

OpenStudy (anonymous):

\[y^l= 2k-y\]

OpenStudy (anonymous):

and for this question k=1=y

OpenStudy (anonymous):

so the answer is \[x^l= 4 \]

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

\[y^l= 2\times 1 - 2 \]

OpenStudy (anonymous):

yah ! it is c

OpenStudy (anonymous):

ok thanks :DD